Construct a 99​% confidence interval to estimate the population mean with x=111 and σ= 32 for the following sample sizes.

​a) n=35

​b) n=44 ​

c) n=65
Answer #1

solution, we have, X = 11 5=32 -0.01 a) consider, n=35 we know, 100(1-2)/- confidence internal for population mean el is giveconsider, n= 65 gg.l. confidence interval for el = (1 + (2:58) 32) → Chi I 10.240) - 1 C 100.76, 121.24)